2 Blanchard Company manufactures a single product that sells for $168 per unit and whose total variable costs are $126 per unit. The companys annual fixed costs are $630,000. Management targets an annual pretax income of $1,050,000. Assume that fixed costs remain at $630,000. (1) Compute the unit sales to earn the target income. Choose Numerator: Choose Denominator:Units to Achieve Target Units to achieve target Skipped (2) Compute the dollar sales to earn the target income Choose Denominator:Dollars to Achieve Target Dollars to achieve target Choose Numerator:
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Selling price per unit                       168.00
Variable cost per unit                     (126.00)
Contribution per unit =168 - 126                         42.00
Contribution margin ratio = 42/168 25.00%
Annual fixed costs               630,000.00
Annual pretax income           1,050,000.00
Desired contribution = 630,000 + 1050,000           1,680,000.00
Unit Sales = 1680,000/42                 40,000.00
Dollar sales=1680,000/25%           6,720,000.00